A cohort study involves a total of 320 participant
Answer #1

Had methane exposure

dermitis

no dermitis

Total

Cumulative

Incidence

Yes

30

120

160

30/160= 0.185

No

10

150

160

10/160 =0.0625

CI Risk Ratio -e

where CIe is the cumulative incidence in the 'exposed' group and CIu is the cumulative incidence in the 'unexposed' group.

0.185 /0.0625 = 2.96

  • A risk ratio > 1 suggests an increased risk of that outcome in the exposed group.